# Gravitation and Spacetime: Emergent from Spinor Interactions - How?

### v2

@@ 1 Introduction
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}+\Lambda g_{\mu\nu}=8\pi GT_{\mu\nu}$ (1)
relate the geometry (left-hand side) to matter content (right-hand side). Quantum mechanics and general relativity become definitely in conflict at length scales as short as the Planck scale, when quantum uncertainty resolves the classical spacetime structure.
@@ 2.3 Curvature Spinors
The Ricci tensor is encoded in the trace-free Ricci spinor $\Phi_{ABA^{\prime}B^{\prime}}$ and the scalar curvature $R$ arises as a spinorial contraction. Physically this relates to the fact, that ultimately all matter is spinorial, fermionic one with spin $\frac{1}{2}$ bosonic one with spin $1$ for the gauge bosons and spin 0 for the Higgs.
@@ 3 Classical Spacetime Geometry in Ashtekar Variables
Ashtekar introduced a reformulation using the self-dual part of the spin connection [9, 10]. With a spatial triad $e^{i}_{a}$ satisfying
@@
where $\Gamma^{a}_{i}$ is the spin connection compatible with $e^{i}_{a}$, and $K^{a}_{i}=K_{ij}e^{aj}$ is the extrinsic curvature in triad form. The factor of $i$ selects the self-dual part.
@@
$\{A^{a}_{i}(x),E^{b}_{j}(y)\}=\delta^{a}_{b}\delta^{i}_{j}\delta^{3}(x,y)$ (21)
@@
$\mathcal{G}_{a}=D_{i}E^{a}_{i}=\partial_{i}E^{a}_{i}+\epsilon_{abc}A^{b}_{i}E^{c}_{i}=0$ (22)
@@
All constraints are polynomial in the fundamental variables. In particular (24) becomes more complicated, if a different e.g. real connection was chosen.
The selfdual connection $A^{a}_{i}$ is complex due to the factor $i$. For real general relativity, reality conditions must be imposed:
@@
Various strategies address this, including a real connection with the Immirzi parameter $\gamma$ [16] and the real Lorentzian formulation [12].
@@ 4 Holonomy and Flux Variables
The momentum conjugate to the holonomy is the flux through a surface $S$:
$E_{S}^{a}[E]=\int_{S}\epsilon^{ijk}E^{a}_{i}dS_{jk}$ (27)
where $n_{k}$ is the normal to $S$.
@@
where the sum is over intersection points $p$ of the path $\gamma$ with surface $S$, with orientations taken into account.
@@ 5.1 Definition of Spin Networks
  • Intertwiners $i_{v}$ at each vertex (invariant tensors coupling the edge representations)
@@
$\Psi_{\Gamma}[A]=\prod_{v\in\gamma}i_{v}\left[\prod_{e\in v}\text{Tr}_{j_{e}}.(h_{e}[A])\right]$ (29)
@@ 5.2 Quantum Geometry of Spin Networks
$\hat{A}(S)\Psi_{\Gamma}=8\pi\gamma l_{P}^{2}\sum_{p\in S\cap\Gamma}\sqrt{j_{p}(j_{p}+1)}\,\Psi_{\Gamma}$ (32)
where $l_{P}=\sqrt{G\hbar/c^{3}}$ is the Planck length and $\gamma$ is the Immirzi parameter. As pointed out in [21], the value of $\gamma$ may be set according to the relevant scale $l_{c}>l_{P}$ for the resulting discrete spatial geometry. In the following section, we propose that, the density of particle interactions is setting the scale $l_{c}$.
@@ 5.3 Surface Entropy from Spin Networks
matching Equation (2), with corrections depending on $\gamma$ and spin statistics.
The microstate counting function $\Omega(A)$ counts the number of distinct spin network configurations that yield a horizon area eigenvalue equal to $A$. Since the area operator has discrete spectrum (Equation (32)), only specific spin network configurations with punctures $\{j_{p}\}$ satisfying
@@
contribute to $\Omega(A)$ with spin values $j_{p}$ at puncture $p$. Puncturing edges correspond to fermions for $j_{p}=\frac{1}{2}$. Hypothetical spin 0 punctures do not yield any area. A single spin $\frac{1}{2}$ puncture yields the minimal area gap $A_{\text{gap}}=4\sqrt{3}\pi\gamma l_{P}^{2}$.
In the canonical LQG approach the Immirzi parameter $\gamma$ mixing connection and extrinsic curvature in the canonical connection is a real number. As discussed in [21], the geometric operators depend on a fundamental length scale $l$ related to the Immirzi parameter via $l=\sqrt{\gamma}l_{P}$. Therefore, with modified $\gamma_{c}$, the relevant critical length $l_{c}$ can be much larger than $l_{P}$, but nevertheless still match the correct black hole entropy with
@@
$S=\frac{k_{B}A}{4l_{P}^{2}}-\frac{3k_{B}}{2}\ln\left(\frac{A}{l_{P}^{2}}\right)+O(1)$ (36)
@@ 6.1 Compact double cones
$i^{+}(\rho)$$\mathcal{J}^{+}(\rho)$$i_{0}(\rho)\cong S^{2}$$\mathcal{J}^{-}(\rho)$$i^{-}(\rho)$$\Sigma_{0}(\rho)$$O$
Figure 3: The double cone $\mathcal{D}(i^{-}(\rho),i^{+}(\rho))$, with future timelike $i^{+}(\rho)$, past timelike $i^{-}(\rho)$, and spatial $i_{0}(\rho)\cong S^{2}$, future and past light cones, $\mathcal{J}^{+}(\rho)$ and $\mathcal{J}^{-}(\rho)$ depicted for a certain value of $\rho>0$. $\Sigma_{0}(\rho)$ denotes the spatial section containing the origin $O$, the causal point of reference.
@@ 6.2 Spin Networks and Particle Spinors within Causal Double Cones
For a given $3+1$ decomposition, we consider a graph $\Gamma\subset\Sigma_{0}\subset\mathcal{D}$ of a spin network inside a spatial hypersurface $\Sigma_{0}$ of a compact double cone $\mathcal{D}$. This framework allows loop quantum gravity to be formulated in a diffeomorphism-invariant algebraic QFT setting, with spin network states accommodating isotony and causality [26].
@@
In this picture, a spinor of representation $j_{e}$ traces a future-directed path in $\mathcal{D}$, which projects to an oriented edge $e$ within $\Gamma$. An interaction event of $n$ spinors at a point $p\in\mathcal{D}$ projects to an intertwiner vertex $v$ of $\Gamma$.
@@
Given this additional interpretation of a spin network, the spinor paths of the standard model particles within $\mathcal{D}$ might at the same time also project to a spin network for the geometry of their space(time) they live in. Even more, space(time) may not need separate spinorial constituents, because the content of particles and their interaction events may be enough to generate the geometry of their spacetime.
Scalar particle fields of spin 0 do not contribute to (area) geometry. The spin 0 field of the SM is the Higgs field. It is supposed to attribute the mass for all massive SM particles, in particular the lepons and quarks, but also the bosons of the weak interaction. The Higgs mass is fixed by
@@
A massive particles provides a Compton clock inside the double cone $\mathcal{D}(p,q)$ along its timelike path between the event $p$, where it is created, and the event $q$, where it is deleted. The massive particles in $\mathcal{D}(p,q)$ then fix the time and length scales of the geometry, breaking its conformal invariance.
@@ 6.3 Conformal Geometry and Spin Networks
$S_{H}^{(c)}=\int\frac{1}{2}e^{(c)}e_{I}^{a(c)}e_{J}^{b(c)}\left(\Omega_{ab}^{IJ(c)}+\frac{1}{\gamma}{}^{*}\Omega_{ab}^{IJ(c)}\right)$ (40)
where $\gamma$ is the Immirzi parameter and $\Omega_{ab}^{IJ(c)}$ is the curvature of the $SL(2,\mathbb{C})$ connection $\omega_{a}^{IJ(c)}$. We follow here the formulation as in [27]. Triads and connection transform conformally as
$e_{a}^{I(c)}$ $=\phi e_{a}^{I}$ (41)
$e^{a(c)}_{I}$ $=\frac{e^{a}_{I}}{\phi}$ (42)
$\omega_{a}^{IJ(c)}$ $=\omega_{a}^{IJ}-2\phi^{-1}\partial_{b}\phi e^{b[I}e_{a}^{J]}$ (43)
and the Poisson bracket of conformal extrinsic curvature and conformally invariant densitized triad is
$\{K_{a}^{i(c)}(x),E_{j}^{b(c)}(y)\}=\delta_{a}^{b}\delta^{i}_{j}\delta^{3}(x,y).$ (44)
@@
$\{K_{a}^{i}(x),E_{j}^{b}\}=\kappa\delta_{a}^{b}\delta^{i}_{j}\delta^{3}(x,y),\quad\{\phi(x),\pi(y)\}=\delta^{3}(x,y)$ (50)
and a conformal constraint
$\phi\pi^{(c)}=\frac{2}{\kappa}K_{a}^{i}E_{i}^{a}+\pi\phi=:\mathcal{S}$ (51)
With real connection variables $A_{a}^{i}=\gamma K_{a}^{i}+\Gamma_{a}^{i}$ the scalar (Hamiltonian) and diffeomorphism constraints are conformally modified to
$H=$ $\frac{\phi^{2}}{2}\epsilon_{i}^{lm}E_{l}^{a}E_{m}^{b}\left[F_{ab}^{i}-\left(\gamma^{2}+(\kappa\phi^{2})^{-2}\right)K_{a}^{j}K_{b}^{k}\epsilon_{jk}^{i}\right]$ (52)
$-E^{ai}E_{i}^{b}\partial_{a}\phi\partial_{b}\phi+2E^{ai}E_{i}^{b}(\nabla_{a}\partial_{b}\phi)\phi,$
$\mathcal{C}_{a}=\frac{1}{\gamma}F_{ab}^{i}E_{i}^{b}+\pi\partial_{a}\phi.$ (53)
@@
$A_{a}^{(c)}=\Gamma_{a}^{i(c)}+\gamma K_{a}^{i(c)}$ (54)
defines a conformal holonomy
$U(A^{(c)},\eta)=P\exp\int A_{a}^{(c)}dy^{a}.$ (55)
Under gauge transformations, conformal holonomies transform as usual via
@@
where $x_{f}^{\eta}$ and $x_{i}^{\eta}$ are the end and starting points of th path $\eta$.
The conformal holonomies yield cylindrical conformal functions as
$\psi_{S}(A^{(c)})=\bigotimes_{l}R^{(j_{l})}(U(A^{(c)},\eta_{l}))\bigotimes_{n}i_{n},$ (57)
where $R^{(j_{l})}$ are the spin $j$ representations of the spinor $SU(2)$ and $i_{n}$ are the $SU(2)$-intertwiners of the representations. (Here we neglect the gauge group $G_{SM}$ of the SM. In a full approach the relevant $G_{SM}$ has to be added consistently on top. For the spin network considered here, bosonic and fermionic spinors are considered equally.) The resulting conformal spin network states live on a scale free topological graph $\Gamma=\cup_{l}\eta_{l}$ within a spatial section $\Sigma_{0}\subset\mathcal{D}$ within a compact causal doublecone $\mathcal{D}$. The edges of the graph represent the particle spinors and the vertices their interactions.
@@
$G$ $=\frac{1}{8\pi\phi_{0}^{2}}$ (58)
and $\phi_{0}=\kappa^{-1/2}$ recovers standard GR coupled to the Standard Model.

### v3

@@ 1 Introduction
$R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi GT_{\mu\nu}$ (1)
relate the pure (without cosmological constant) geometry (left-hand side) to matter (outside vacuum) content (right-hand side). Quantum mechanics and general relativity become definitely in conflict at length scales as short as the Planck scale, when quantum uncertainty resolves the classical spacetime structure.
@@ 2.3 Curvature Spinors
Likewise, the trace-free Ricci tensor is encoded in a trace-free Ricci spinor $\Phi_{ABA^{\prime}B^{\prime}}$ and the scalar curvature $R$ arises as a spinorial contraction. Physically this relates to the fact, that ultimately all matter is spinorial, fermionic one with spin $\frac{1}{2}$ bosonic one with spin $1$ for the gauge bosons and spin 0 for the Higgs.
@@ 3 Classical Spacetime Geometry in Ashtekar Variables
Ashtekar introduced a reformulation using the self-dual part of the spin connection [9, 10]. With a spatial triad $e^{i}_{a}$ ($a$ is now an internal $SU(2)$ index) on hypersurface $\Sigma_{t}$, satisfying
@@
where $\Gamma^{a}_{i}$ is the spin connection compatible with $e^{a}_{i}$, and $K^{a}_{i}=K_{ij}e^{aj}$ is the extrinsic curvature in triad form. The factor of $i$ selects the self-dual part.
@@
$\{A^{a}_{i}(x),E^{j}_{b}(y)\}=\delta^{a}_{b}\delta^{j}_{i}\delta^{3}(x,y)$ (21)
@@
$\mathcal{G}_{a}=D_{i}E^{i}_{a}=\partial_{i}E^{i}_{a}+\epsilon_{ab}^{c}A^{b}_{i}E^{i}_{c}=0$ (22)
@@
All constraints are polynomial in the fundamental variables. In particular (24) becomes more complicated, if a different e.g. real connection was chosen. The selfdual connection $A^{a}_{i}$ is complex:
@@
For real general relativity then, reality conditions must be imposed. Various strategies address this, including a real connection with the Immirzi parameter $\gamma$ [16] and the real Lorentzian formulation [12].
@@ 4 Holonomy and Flux Variables
A conjugate variable to the holonomy is the flux through a surface $S$, intersecting $\gamma$ transversally,
$E_{S}^{a}[E]=\int_{S}\epsilon^{ijk}E^{a}_{i}dS_{jk}.$ (27)
@@
where the sum is over intersection points $p$ of the path $\gamma$ with surface $S$ at its segment $\gamma_{p}$, with orientations taken into account.
@@ 5.1 Definition of Spin Networks
  • Intertwiners $i_{v}$ at each vertex (i.e. invariant recouplings of the spin representations $R^{(j_{e})}$ of the oriented incoming and outgoing edges at each vertex)
@@
$\Psi_{\Gamma}[A]=\prod_{v\in\gamma}i_{v}\left[\prod_{e\in v}R^{(j_{e})}(h_{e}[A])\right]$ (29)
@@ 5.2 Quantum Geometry of Spin Networks
$\hat{A}(S)\Psi_{\Gamma}=8\pi\gamma l_{c}^{2}\sum_{p\in S\cap\Gamma}\sqrt{j_{p}(j_{p}+1)}\,\Psi_{\Gamma}$ (32)
where the common standard choice for the critical length scale is $l_{c}=l_{P}=\sqrt{G\hbar/c^{3}}$, the Planck length, and $\gamma$ is the Immirzi parameter. As pointed out in [21], the value of $\gamma$ may be set according to the relevant scale $l_{c}>l_{P}$ for the resulting discrete spatial geometry. In the following section, we propose that, the density of particle interactions is setting the scale $l_{c}$.
@@ 5.3 Surface Entropy from Spin Networks
matching Equation (2), with corrections depending on $\gamma$ and spin statistics. The microstate counting function $\Omega(A)$ counts the number of distinct spin network configurations that yield a horizon area eigenvalue equal to $A$. Since the area operator has discrete spectrum (Equation (32)), only specific spin network configurations with punctures $\{j_{p}\}$ satisfying
@@
contribute to $\Omega(A)$ with spin values $j_{p}$ at puncture $p$. Puncturing edges with $j_{p}=\frac{1}{2}$ orrespond to fermions. Hypothetical spin 0 punctures do not yield any area. A single spin $\frac{1}{2}$ puncture yields the minimal area gap $A_{\text{gap}}=4\sqrt{3}\pi\gamma l_{P}^{2}$.
In the canonical LQG approach the Immirzi parameter $\gamma$, mixing connection and extrinsic curvature into the canonical connection, is a real number. As discussed in [21], the geometric operators depend on a fundamental length scale $\ell$ related to the Immirzi parameter via $\ell=\sqrt{\gamma}l_{P}$. Therefore, with modified $\gamma_{c}$, the relevant critical length $l_{c}$ can be much larger than $l_{P}$, but nevertheless still match the correct black hole entropy with
@@
$\frac{S}{k_{B}}=\frac{1}{4}\frac{A}{l_{P}^{2}}-\frac{1}{2}\ln\left(\frac{A}{l_{P}^{2}}\right)+O(1)$ (36)
@@ 6.1 Compact double cones
$i^{+}(\rho)$$\mathcal{J}^{+}(\rho)$$i^{0}(\rho)\cong S^{2}$$\mathcal{J}^{-}(\rho)$$i^{-}(\rho)$$\Sigma_{0}(\rho)$$O$
Figure 3: The double cone $\mathcal{D}(i^{-}(\rho),i^{+}(\rho))$, with future timelike $i^{+}(\rho)$, past timelike $i^{-}(\rho)$, and spatial $i^{0}(\rho)\cong S^{2}$, future light cone $\mathcal{J}^{-}(\rho)$ of $i^{-}(\rho)$ and past light cone $\mathcal{J}^{+}(\rho)$ of $i^{+}(\rho)$, depicted for a certain value of $\rho>0$. $\Sigma_{0}(\rho)$ denotes the spatial section containing the origin $O$, the causal point of reference.
@@ 6.2 Spin Networks and Particle Spinors within Causal Double Cones
For a given $3+1$ decomposition, we consider a spin network $\Gamma$ for a graph $\gamma\subset\Sigma_{0}\subset\mathcal{D}$ of a spin network inside a spatial hypersurface $\Sigma_{0}$ of a compact double cone $\mathcal{D}$. This framework allows loop quantum gravity to be formulated in a diffeomorphism-invariant algebraic QFT setting, with spin network states accommodating isotony and causality [26].
@@
In this picture, a spinor of representation $j_{e}$ traces a future-directed path in $\mathcal{D}$, which projects to an oriented edge $e$ within $\Gamma$. An interaction event of $n$ spinors at a point $p\in\mathcal{D}$ projects to an intertwiner vertex $v$ of $\Gamma$. Each intertwiner $i_{v}$ at a $v$ encodes the entanglement of the $n$ spinors.
@@
Given this additional interpretation of a spin network, the spinor paths of the standard model particles within $\mathcal{D}$ might at the same time also project to a spin network for the geometry of their space(time) they live in. Even more, space(time) may not need separate spinorial constituents, because the content of particles and their interaction events may be enough to generate the geometry of their space(time).
Scalar particle fields of spin 0 do not contribute to (area) geometry. The spin 0 field of the SM is the Higgs field. It is supposed to attribute the mass for all massive SM particles, in particular the lepo ns and quarks, but also the bosons of the weak interaction. The Higgs mass is fixed by a Higgs field vacuum expectation
@@
A massive particle provides a Compton clock inside the double cone $\mathcal{D}(p,q)$ along its timelike path between the event $p$, where it is created, and the event $q$, where it is deleted. The massive particles in $\mathcal{D}(p,q)$ then fix the time and length scales of the geometry, breaking its conformal invariance.
@@ 6.3 Conformal Geometry and Spin Networks
$S_{H}^{(c)}=\int\frac{1}{2}e^{(c)}e_{a}^{\mu(c)}e_{b}^{\nu(c)}\left(\Omega_{\mu\nu}^{ab(c)}+\frac{1}{\gamma}{}^{*}\Omega_{\mu\nu}^{ab(c)}\right)$ (40)
where $\gamma$ is the Immirzi parameter, Latin indices $a,b,\ldots$ label the representation basis for the gauge group $SL(2,\mathbb{C})$, and $\Omega_{\mu\nu}^{ab(c)}$ is the curvature of the $SL(2,\mathbb{C})$ connection $\omega_{\mu}^{ab(c)}$. We follow here the formulation as in [27]. After $3+1$ decomposition, we have triads and connections transforming conformally as
$e_{i}^{a(c)}$ $=\phi e_{i}^{a}$ (41)
$e^{i(c)}_{a}$ $=\frac{e^{i}_{a}}{\phi}$ (42)
$\omega_{i}^{ab(c)}$ $=\omega_{i}^{ab}-2\phi^{-1}\partial_{j}\phi e^{j[a}e^{b]}_{i}$ (43)
with spatial Latin indices $i,j,\ldots$. The Poisson bracket of conformal extrinsic curvature and conformally invariant densitized triad is
$\{K_{i}^{a(c)}(x),E_{b}^{j(c)}(y)\}=\delta_{i}^{j}\delta^{a}_{b}\delta^{3}(x,y).$ (44)
@@
$\{K_{i}^{a}(x),E_{b}^{j}\}=\kappa\delta_{a}^{b}\delta^{j}_{i}\delta^{3}(x,y),\quad\{\phi(x),\pi(y)\}=\delta^{3}(x,y)$ (50)
and a conformal constraint quantity
$\phi\pi^{(c)}=+\pi\phi+\frac{2}{\kappa}K_{i}^{a}E_{a}^{i}$ (51)
With real connection variables $A_{i}^{a}=\gamma K_{i}^{a}+\Gamma_{i}^{a}$ the scalar (Hamiltonian) and diffeomorphism constraints are conformally modified to
$H=$ $\frac{\phi^{2}}{2}\epsilon_{a}^{bc}E_{l}^{i}E_{c}^{j}\left[F_{ij}^{a}-\left(\gamma^{2}+1)(\kappa\phi^{2})^{-2}\right)K_{i}^{b}K_{j}^{c}\epsilon_{bc}^{a}\right]$ (52)
$+E^{ia}E_{a}^{j}[-\partial_{i}\phi\partial_{j}\phi+2(\nabla_{i}\partial_{j}\phi)\phi],$
$\mathcal{H}_{i}=\frac{1}{\gamma}F_{ij}^{a}E_{a}^{j}+\pi\partial_{i}\phi.$ (53)
@@
$A_{i}^{a(c)}=\Gamma_{i}^{a(c)}+\gamma K_{i}^{a(c)}\quad,\quad A_{i}^{(c)}=A_{i}^{a(c)}\tau_{a}$ (54)
defines a conformal holonomy along an oriented path $\eta$ as
$h_{\eta}[A^{(c)}]=U(A^{(c)},\eta)=P\exp\int_{\eta}A_{i}^{(c)}dx^{i}\in SU(2).$ (55)
Under $SU(2)$ gauge transformations, conformal holonomies transform as usual via
@@
where $x_{f}^{\eta}$ and $x_{i}^{\eta}$ are the end and starting points of the path $\eta$.
The conformal holonomies yield cylindrical conformal functions related to $\Gamma=(\gamma,j_{e},i_{v})$ as
$\Psi_{\Gamma}[A^{(c)}]=\prod_{v\in\gamma}i_{v}\left[\prod_{e\in v}R^{(j_{e})}(h_{e}[A^{(c)}])\right],$ (57)
where $R^{(j)}$ is the spin $j$ representation of the gauge group $SU(2)$ and $i_{v}$ is the $SU(2)$-intertwiner of the representations adjacent to vertex $v\in\gamma$. (Here we neglect the gauge group $G_{SM}$ of the SM. In a full approach the relevant $G_{SM}$ has to be added consistently on top. For the spin network considered here, bosonic and fermionic spinors are considered equally.) The resulting conformal spin network states live on a scale free topological graph $\gamma$ within a spatial projection section $\Sigma_{0}\subset\mathcal{D}$ within a compact causal doublecone $\mathcal{D}$. The edges of the graph represent the particle spinors and the vertices their interaction entanglement.
@@
$G$ $=\frac{1}{8\pi\phi_{0}^{2}}\quad,$ (58)
i.e. $\phi_{0}=\kappa^{-1/2}$ recovers standard GR as coupled to the Standard Model.